Show that there is no real number, k for which the equation, x4 - 3x2 + k =0 has two distinct roots in the interval [2,3].


1
Expert's answer
2021-06-01T12:39:37-0400

Given

x4 - 3x+ k =0


we want to find value of k, for which the equation, x4 - 3x+ k =0 has two distinct roots in the interval [2,3].

we write equation in this form

"(x^2)^2-3(x^2)+k=0"

let two distinct root is x1 and x2

according to question


"2\leq x_1, x_2 \leq 3\\
4\leq x_1^2, x_2^2 \leq 9\\"

we know ,

"x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}
\\"

hence

"4\leq \frac{3\pm\sqrt{9-4k}}{2} \leq 9\\
8\leq {3\pm\sqrt{9-4k}}{} \leq 18\\
5\leq \pm\sqrt{9-4k}\leq 15.......(1)
\\
now \space take \space + or- \space one by one\\
5\leq \sqrt{9-4k}\leq 15....(2)\\
and\\

5\leq -\sqrt{9-4k}\leq 15\\
or\\
-5\geq \sqrt{9-4k}\geq -15.....(3)\\
from(2) and(3)\\
there\space no\space existing\space k \space which \space \\satisfied \space eq(2)and(3) 
\\"

hence we say that

there is no real number, k for which the equation, x4 - 3x+ k =0 has two distinct roots in the interval [2,3].